OpenStudy (anonymous):

Find a quadratic function to model the values in the table. Predict the value of y for x = 4. X|Y -1|-3 0|-2 3|-35

OpenStudy (anonymous):

the constant must be \(-2\) so you have \[y=ax^2+bx-2\] as a start now we can set up a couple equations in \(a\) and \(b\)

OpenStudy (anonymous):

let \(x=-1\) you get \[a-b-2=-3\] or\[a-b=-1\]

OpenStudy (anonymous):

let \(x=3\) you get \[9a+3b-2=-35\] or \[9a+3b=-33\]

OpenStudy (anonymous):

solve the system \[a-b=-1\]\[9a+3b=-33\] to get your answer

OpenStudy (anonymous):

multiply first equation by 3, add it to the second get \[12a=-36\] so \(a=-3\) therefore \(b=-2\) and you get \[y=-3x^2-2b-2\]

OpenStudy (anonymous):

oops \[y=-3a^2-2x-2\]

OpenStudy (anonymous):

so it would be y=-3a^2-2x-2;-58

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

yes for the first one

OpenStudy (anonymous):

it would be \[y=-3x^2-2x-2\] with an \(x\) not an \(a\)

OpenStudy (anonymous):

thanks so much. i've been working on this thing all weekend....

OpenStudy (anonymous):

wait so 4 is going to be -56?

OpenStudy (anonymous):

@satellite73